OpenStudy (anonymous):

Finish these three solutions to this equation. y=x-1/2 (___, 1/2 ), (___, 3/2 ), (___, 5/2 )

OpenStudy (anonymous):

1/2=x-1/2 3/2=x-1/2 5/2=x-1/2 i think you can do this ;)

OpenStudy (31356):

This is the first one -1=1/2

OpenStudy (31356):

Medal?

OpenStudy (31356):

Solve it by a graphing calculator, Anna, it will help you a ton better!